1) find the power of w(t)

2) find the complex fourier series of coefficient of w(t)2. (10 points) Consider the waveform shown below: w(t) 2 w t -4 -2 0 2
Answer #1

page-1 * Given wilt) is a periodic signal. ak(t) 2 (0,2) >t 2 4 6 -4 (2,0) (0, 2) (20) (y-y) = X-X, (WLt) -2) = 0-2 (t-o) 2-0é jant- page-4 1+1 Ž T? JAT -H h 2 jnt I 1 Cn = INT inwot Cn X(t) = 2 n =-00 j wo -T Cn = int intt 1 alt) = int N=-00